# Review-based Recommender Systems: A Survey of Approaches, Challenges and Future Perspectives

### v2

@@ 9. DISCUSSION AND FUTURE DIRECTIONS
The study of review-based recommender systems highlights both the progress and challenges in this area. Despite encountering several challenges, review-based recommender systems offer a significant advancement in recommendation technology by leveraging the fine-grained insights present in user-generated reviews to provide highly personalized suggestions. These systems explore the textual content of reviews to grasp users’ sentiments, preferences, and fine-grained opinions about products or services. This deeper analysis facilitates enhanced personalization, as the systems can comprehend individual tastes and requirements with greater precision, resulting in recommendations that closely align with each user’s specific preferences. Moreover, review-based systems improve accuracy by considering qualitative information present in reviews, such as detailed explanations, experiences, and comparisons, which may not be captured by ratings.
Additionally, review-based recommender systems excel at capturing detailed information about products or services, providing recommendations that consider specific attributes or characteristics that matter most to users. This adaptability to diverse preferences ensures that the recommendations can meet various user interests, accommodating different priorities such as quality, price, or specific features. By leveraging user-generated content, review-based systems offer a trustworthy recommendation process, fostering confidence in users’ decision-making. However, these improvements also bring up important questions about ethics, how transparent the algorithms are, and the importance of maintaining user trust and data security.
Given the complexities and rapid advancements in Review-based Recommender System, the following subsection outlines several promising areas for future research in this field:
@@
9.1. Enhancing Interpretability and Trust
Future advancements should focus on the interpretability and trust. Users benefit not only from accurate recommendations but also from understanding why specific items are recommended. Enhancing interpretability involves developing models that can explicitly link recommendations to specific aspects mentioned in reviews, such as quality, usability, or aesthetics. This can be achieved through more sophisticated aspect-based analysis and by incorporating explanation modules that can convey the rationale behind recommendations in an easily understandable manner. Building systems that offer transparent reasoning behind their suggestions could significantly improve user trust and satisfaction.
9.2. Leveraging Multimodal Data
As user reviews increasingly include multimedia content in reviews, such as images, videos, and audio clips, future recommendation systems could benefit from a multimodal approach that integrates these diverse types of data. This involves not only analyzing textual content but also understanding and extracting sentiments and preferences from images and videos. For instance, visual sentiment analysis from product images or user-generated content in reviews could offer additional insights into user preferences, enhancing the recommendation process. Research in multimodal learning and its integration into recommender systems remains an under-exploited area with considerable potential for innovation.
9.3. Leveraging Multi-criteria Rating Data
Incorporating multi-criteria ratings, which reflect specific aspects of user satisfaction, can enhance both the performance and interpretability of Review-based recommender systems. Tripadvisor  [^11] https://www.cs.virginia.edu/~hw5x/Data/LARA/TripAdvisor/, for instance, collects ratings on different criteria including location, Service, Cleanliness, Value, Sleep Quality, etc. The ratings on the individual criteria represent users’ fine-grained preferences because not all the criteria are equally important to a user. Therefore, integrating these criteria rating information into the review and overall rating features could potentially improve the performance and provide better interpretability. Furthermore, multi-criteria rating information is valuable in mitigating the issue of data sparsity. In cases where reviews are sparse, leveraging multiple criteria can aid in learning user preferences. Recently, MRRRec (Hasan et al., 2022), and (Hasan and Ding, 2023) demonstrated the effectiveness of integrating multicriteria rating data with review to enhance the performance of the recommender system. However, a sophisticated approach is required to take full advantage of the utilizing criteri-rating information into reviews.
9.4. Addressing Data Sparsity
@@ 9.5. Ethical Considerations and Bias Mitigation
As recommender systems increasingly influence what information and products users are exposed to, ethical considerations become paramount. Future research must prioritize the development of mechanisms to identify and mitigate the biases in review-based recommender systems. This includes biases introduced by demographic factors, manipulation through fake reviews, or systemic biases inherent in the data collection process. Ensuring fairness, diversity, and representativeness in recommendations is essential for building equitable systems that serve a broad user base effectively.
@@ 9.7. Cross-Domain Recommendations
Exploring ways to apply knowledge from one product area to another and creating a common system to understand user likes and item features are key steps to make recommendation systems better and more flexible across different types of products. Domain adaptation techniques help by using what the system has learned in one area to improve recommendations in another, finding similarities between them (Khan et al., 2017). At the same time, creating a single framework that captures what users like and item details across all areas can help make recommendations more consistent and relevant, no matter the product type. These methods could lead to better recommendations across various product categories, making suggestions more personalized and useful to users.
@@ 9.8. Large Language Models and Recommendations
Large Language Models (LLMs) can play a crucial role in generating recommendations by leveraging their natural language understanding capabilities to analyze customer reviews and generate insightful, relevant suggestions (Fan et al., 2023; Laskar et al., 2023a, b). Recent studies have explored the potential of LLMs in recommender systems, with a focus on prompting engineering (Xu et al., 2024; He et al., 2023). These studies have highlighted the effectiveness of LLMs in recommendation tasks, particularly in understanding context, learning user preferences, and generating relevant recommendations. In developing review-based recommendations, LLMs can facilitate sentiment analysis, key theme extraction, personalized recommendations, predictive analysis, and review summarization through methods like zero-shot learning (Kojima et al., 2022), demonstrations (Dang et al., 2022), or fine-tuning (Raza and Ding, 2023). This strategy capitalizes on the advanced natural language understanding capabilities of LLMs, offering customers insightful, targeted recommendations that enhance decision-making and satisfaction (Liu et al., 2023).
@@ 9.9. Responsible Recommendations
Most research on recommender systems has traditionally aimed at making the recommended results more accurate. However, the importance of ethical considerations and the impact of these systems on society has increasingly come into focus (Wang et al., 2022). There is a need for focusing on the aspects of privacy, diversity, and fairness, and how they interconnect with the notion of responsibility in aspect-based recommender systems (Elahi et al., 2022). While, the current literature often lacks a systematic exploration of these integrated approaches, particularly in the context of review-based recommender systems, these systems could greatly benefit from such responsible practices.
@@ 10. CONCLUSIONS AND FUTURE WORK
This paper presents a comprehensive exploration of review-based recommender systems, demonstrating their significant potential. We introduce various classification schemes based on feature extraction and review integration strategies. Through a detailed analysis of state-of-the-art methodologies, we show how leveraging user-generated reviews can substantially enhance the personalization and accuracy of recommendations, moving beyond the limitations of traditional rating-based systems. We identify the key challenges, including the complexity of learning user/item representation from review, review-integration strategies, the need for scalable solutions to handle the vast datasets, and the importance of ethical considerations in data handling. Despite these challenges, advancements in machine learning, deep learning, and NLP techniques offer promising avenues for extracting valuable insights from reviews, thereby enriching recommender systems with detailed understandings of user preferences.
Looking forward, the integration of multimodal data presents an exciting opportunity for further enriching review-based recommender systems, enabling a holistic approach to understanding user sentiments and preferences. Addressing data sparsity and the cold start problem remains a critical area for future research. Moreover, as recommender systems increasingly influence user choices and experiences, ethical considerations around bias mitigation and data privacy must be central to future developments. Ensuring that recommender systems are fair, and transparent, and respect user privacy will be paramount in fostering trust and acceptance among users.
@@ 11. Acknowlegements
This work is supported by the Natural Sciences and Engineering Research Council of Canada and the York Research Chairs program. Jimmy Huang (jhuang@yorku.ca) and Chen Ding (cding@torontomu.ca) are the contact authors of this paper. We would like to extend our gratitude to Syed Faizan for his invaluable assistance in gathering and organizing several papers

### v3

@@ 10. DISCUSSION AND FUTURE DIRECTIONS
Despite challenges, review-based recommender systems advance the recommendation process by using fine-grained insights from user reviews to deliver personalized suggestions. These systems explore the textual content of reviews to grasp users’ sentiments, preferences, and fine-grained opinions about products or services. This deeper analysis facilitates enhanced personalization, as the systems can comprehend individual tastes and requirements with greater precision, resulting in recommendations that closely align with each user’s specific preferences. Moreover, review-based systems improve accuracy by considering qualitative information present in reviews, such as detailed explanations, experiences, and comparisons, which may not be captured by ratings.
Additionally, review-based recommender systems excel at capturing detailed information about products or services, providing recommendations that consider specific attributes or characteristics. This adaptability to diverse preferences ensures that the recommendations can meet various user interests, accommodating different priorities such as quality, price, or specific features. By leveraging user-generated content, review-based systems offer a trustworthy recommendation process, fostering confidence in users’ decision-making.
10.1. Enhancing Interpretability and Trust
Future advancements should focus on interpretability and trust. Users benefit not only from accurate recommendations but also from understanding why specific items are recommended. Enhancing interpretability involves developing models that can explicitly link recommendations to specific aspects mentioned in reviews, such as quality, usability, or aesthetics. This can be achieved through more sophisticated aspect-based analysis and by incorporating explanation modules that can convey the rationale behind recommendations in an easily understandable manner. Building systems that offer transparent reasoning behind their suggestions could significantly improve user trust and satisfaction.
10.2. Leveraging Multi-modal Data
As user reviews increasingly include multimedia content in reviews, such as images, videos, and audio clips, future recommendation systems could benefit from a multi-modal approach that integrates these diverse types of data. This involves not only analyzing textual content but also understanding and extracting sentiments and preferences from images and videos. For instance, visual sentiment analysis from product images or user-generated content in reviews could offer additional insights into user preferences, enhancing the recommendation process. Research in multi-modal learning and its integration into recommender systems remains an under-exploited area with considerable potential for innovation.
10.3. Leveraging Multi-criteria Rating Data
Incorporating multi-criteria ratings, which reflect specific aspects of user satisfaction, can enhance both the performance and interpretability of review-based recommender systems. Tripadvisor [^9] https://www.cs.virginia.edu/~hw5x/Data/LARA/TripAdvisor/, for instance, collects ratings on different criteria including location, Service, Cleanliness, Value, Sleep Quality, etc. The ratings on the individual criteria represent users’ fine-grained preferences because not all the criteria are equally important to a user. Therefore, integrating these criteria ratings into the review and overall ratings could potentially improve the performance and provide better interpretability. Furthermore, multi-criteria rating information is valuable in mitigating the issue of data sparsity. In cases where reviews are sparse, leveraging multiple criteria can aid in learning user preferences. Recently, MRRRec (Hasan et al., 2022), and (Hasan and Ding, 2023) demonstrate the effectiveness of integrating multi-criteria rating data with reviews to enhance the performance of the recommender system. However, a sophisticated approach is required to take full advantage of the utilizing criteria ratings into reviews.
10.4. Addressing Data Sparsity
@@ 10.5. Ethical Considerations and Bias Mitigation
As recommender systems increasingly influence what information and products users are exposed to, ethical considerations become paramount. Future research must prioritize the development of mechanisms to identify and mitigate the biases in review-based recommender systems. This includes biases introduced by demographic factors, manipulation through fake reviews, or systemic biases inherent in the data collection process. Ensuring fairness, diversity, and representativeness in recommendations is essential for building equitable systems that serve a broad user base effectively. Review-based recommender systems can benefit from including transparency, ethics, user trust, and data security for fairness and reliability.
@@ 10.7. Cross-Domain Recommendations
Exploring ways to apply knowledge from one product area to another and creating a common system to understand user likes and item features are key steps to make recommendation systems better and more flexible across different types of products. Domain adaptation techniques help by using what the system has learned in one area to improve recommendations in another, finding similarities between them (Khan et al., 2017). At the same time, creating a single framework that captures what users like and item details across all areas can help make recommendations more consistent and relevant, no matter the product type. These methods could lead to better recommendations across various product categories, making suggestions more personalized and useful.
@@ 10.8. Large Language Models and Recommendations
LLMs can play a crucial role in generating recommendations by leveraging their natural language understanding capabilities to analyze customer reviews and generate insightful, relevant suggestions (Fan et al., 2023; Laskar et al., 2023a, b). Recent studies have explored the potential of LLMs in recommender systems, with a focus on prompting engineering (Xu et al., 2024; He et al., 2023). These studies have highlighted the effectiveness of LLMs in recommendation tasks, particularly in understanding context, learning user preferences, and generating relevant recommendations. In developing review-based recommendations, LLMs can facilitate sentiment analysis, key theme extraction, personalized recommendations, predictive analysis, and review summarization through methods like zero-shot learning (Kojima et al., 2022), demonstrations (Dang et al., 2022), or fine-tuning (Raza and Ding, 2023). Additionally, LLMs can improve the processing of unstructured data, such as handling informal language, emojis, and unique formatting used in reviews, which traditional models may struggle to interpret accurately. By handling these complexities more easily, LLMs can provide a more accurate understanding of customer feedback and sentiment, enhancing the overall recommendation quality (Liu et al., 2023b). This strategy capitalizes on the advanced natural language understanding capabilities of LLMs, offering customers insightful, targeted recommendations that enhance decision-making and satisfaction (Liu et al., 2023b).
@@ 10.9. Responsible Recommendations
Most research on recommender systems has traditionally aimed at making the recommended results more accurate. However, the importance of ethical considerations and the impact of these systems on society have increasingly come into focus (Wang et al., 2022). There is a need for focusing on the aspects of privacy, diversity, and fairness, and how they interconnect with the notion of responsibility in aspect-based recommender systems (Elahi et al., 2022). While the current literature often lacks a systematic exploration of these integrated approaches, particularly in the context of review-based recommender systems, these systems could greatly benefit from such responsible practices.
@@ 11. CONCLUSIONS AND FUTURE WORK
This paper presents a comprehensive exploration of review-based recommender systems. We classify these systems using two frameworks: (a) review utilization strategies (generic review methods, aspect-based methods, rating-review fusion, and ratings and aspects fusion) and (b) methodologies (probabilistic models, deep learning, and miscellaneous models). Through a detailed analysis of state-of-the-art methodologies, we show how leveraging user-generated reviews can enhance the personalization and accuracy of recommendations, surpassing the limitations of traditional rating-based systems. We identify the key challenges, including the complexity of learning user/item representations from reviews, review integration strategies, the need for scalable solutions to handle vast datasets, and the importance of ethical considerations in data handling. Despite these challenges, advancements in machine learning, deep learning, and NLP techniques offer promising avenues for extracting valuable insights from reviews, thereby enriching recommender systems with detailed understandings of user preferences.
Looking forward, the integration of multi-modal data presents an exciting opportunity for further enriching review-based recommender systems, enabling a holistic approach to understanding user sentiments and preferences. Addressing data sparsity and the cold start problem remains a critical area for future research. Moreover, as recommender systems increasingly influence user choices and experiences, ethical considerations around bias mitigation and data privacy must be central to future developments. Ensuring these systems are fair, transparent, and privacy-compliant will be paramount in fostering long-term trust and acceptance among users.
@@ 12. Acknowlegements
This work is supported by the Natural Sciences and Engineering Research Council of Canada and the York Research Chairs program. Jimmy Huang (jhuang@yorku.ca) and Chen Ding (cding@torontomu.ca) are the contact authors of this paper. We would like to extend our gratitude to Syed Faizan for his invaluable assistance in gathering and organizing several papers. We sincerely thank the associate editor and all the anonymous reviewers for their comprehensive feedback and thoughtful insights.